-- Main.Walth - 19 Jan 2013

5:33pm: Changing masks, patchy clouds

6:03pm: Starting observation GUIs for the night

Decision to use the GG495 blocking filter for the observations tonight, there was concern over 2nd order contamination in the red side.

A773 flats were not taken from last night, glich perhaps? They need to be taken without the blocking filter.

6:41pm: Starting standard G191-B2B acquisition, clearer, but still patchy clouds

6:49pm: Starting standard G191-B2B integration, problem with shell ripping, delay

7:09pm: MACS0647 starting acquisition, problem changing the mask, same as previous night.

7:28pm: Acquisition failed again, seems to be related to elevation, elevation 42deg, 1.5 airmass.

7:36pm: Gave up on mask, doing imaging for MACS0647 until we can put the mask in

7:48pm: Lost guide star, switching back to spectroscopy, starting MACS0647 acquisition, mask changing failed.

We re-tried a few times, and then it started to work.

8:05pm: Guider picked wrong guide star, huge offset in the alignment.

8:18pm: MACS0647 start science, seeing ~1"

10:00pm: A611 acquisition

10:13pm: A611 start science, seeing 0.65"

12:23am: MACS0717 acquisition

12:38am: MACS0717 start science, seeing 0.6"

3:55am: Feige 34 standard acquisition

4:01am: Feige 34 integration, seeing 0.73"

4:14am: MACS1149 acquisition

4:32am: MACS1149 start science, seeing 0.68"

6:53am: Twilight flats for all masks, seems clear out

7:17am: Starting calibrations
Topic revision: r5 - 19 Jan 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback